zarkam21:

a vector has a y-component with length 16 and an x-component with length 14. At what angle is the vector?

zarkam21:

@Vocaloid

Vocaloid:



Vocaloid:

if you have the adjacent & opposite sides what trig ratio do we use?

zarkam21:

tan

Vocaloid:

good so tan(theta) = 16/14 to solve for theta, take the arctan of both sides

Vocaloid:

theta = arctan(16/14)

zarkam21:

wait so is this for the x component?

zarkam21:

?

Vocaloid:

we already know the x and y components we are solving for theta

zarkam21:

oh I got confused

zarkam21:

0.86

Vocaloid:

that's radians (which is technically fine but your teacher is probably expecting degrees)

Vocaloid:

so 48.81 degrees = your answer

zarkam21:

and how did we get that

Vocaloid:

degrees = (180/pi)*radians or just set your calculator to degree mode

zarkam21:



zarkam21:

what is the angle theta in the triangle?

Vocaloid:

same method as before, use tan

zarkam21:

9/4.2

Vocaloid:

good, tan(theta) = 9/4.2 then solve for theta

zarkam21:

4/9

Vocaloid:

tan(theta) = 9/4.2 take the arctan of both sides

zarkam21:

2.14

zarkam21:

?

Vocaloid:

arctan(9/4.2) = ?

zarkam21:

1.13

Vocaloid:

good (you can convert that to degrees if you'd like)

zarkam21:

so 1.13 degrees

Vocaloid:

it's 1.13 radians, convert that to degrees if you wish

zarkam21:

65

Vocaloid:

good, that's it